OpenStudy (anonymous):

log 5 x ^2 + log5 x^3 =0

OpenStudy (anonymous):

\[\log_{5 } x)^2 + \log_{5} x^3 = 0\]

OpenStudy (calchelp):

\[\log_{5} (x^2 * x^3) = 0 \]\[(x^5) = 5^0\]\[x^5 = 1\]\[x = 1.\] I think that is the answer